Jon is three times as old as Cal, while Ron is 12 years older that Cal. Eight years from now. the sum of their ages will be 81. How old are Jon,Cal and Ron ?
Found 2 solutions by ikleyn, josgarithmetic:
Answer by ikleyn(52807)

.

x for Cal, 3x for Jon, and (x+12) for Ron.


Write equation for their total age now, which is 81 - 3*8 = 57

    x + 3x + (x+12) = 57


Simplify and find x

    5x = 57 - 12 = 45

    x            = 45/5 = 9.


ANSWER.  Cal is 9 years old;  Jon is 3*9 = 27 years old; Ron is 9+12 = 21 years old.

Solved.



Answer by josgarithmetic(39620)

          NOW            8 YEARS FUTURE

JON        3c               3c+8

CAL         c                c+8

RON        c+12              c+20

Sum                          81